It is observed that the number of bacteria present in the portion of the small intestine of an animal grows exponentially. Initially,there are 10,000 bacteria present. Three hours later,the bacteria grow to 500,000. Assuming no bacteria due in the progress,how many bacteria will there be after one day? After how many hours will the bacteria double it's number?


1
Expert's answer
2021-06-22T18:03:53-0400

solution

10,000er(3) =500,000


10,000er(3)/10000 =500,000/10,000

introduce (ln) function

ln e3r =50

3r=ln 50


3r/3=ln 50/3 use a calculator

r=1.304007668

exponential function becomes 10,000e1.304007668(t)


Then the population after 1 day becomes

1day= 24 hours

Therefore

10000e1.304007668(24)=(3.906249948 x 10) 17


b)how many hours the bacteria will take to double its number


10,000e1.304007668t=20,000

10,000e1.304007668t/10,000=20000/10000

ln e 1.304007668t=ln 2


1.304007668t/1.304007668=ln 2/1.304007668

t= ln2/1.304007668

t=0.5 hours or 30 minutes